OpenStudy (anonymous):

You work at a T-shirt printing business. Of the 2,800 T-shirts shipped, 396 have a defect. What is the experimental probability that a T-shirt has a defect? Write your answer as a percent, to the nearest tenth of a percent

OpenStudy (anonymous):

The experimental probability is the number that have a defect over the total number then multiply what you get by 100. so 396/2800 times 100

OpenStudy (anonymous):

so its \[\frac{396}{2800} \times 100\] because its \[\frac{defects}{total} = 0.1414\] now multiply by 100 to get the percentage =14.1%